OpenStudy (anonymous):

Integral sec t (tan t)^2 dt

OpenStudy (anonymous):

u = sect

OpenStudy (turingtest):

hah, you got that from that sqrt(x-2)/sqrt(x-1) thing, right?

OpenStudy (turingtest):

sect(tan^2t)=sect(sec^2t-1)=sec^3t-sect

OpenStudy (anonymous):

I did

OpenStudy (turingtest):

\[\int\sec^3 tdt\]\[u=\sec t\]\[dv=\sec^2 t\]integrate by parts

OpenStudy (anonymous):

I've gotten that far in the equation already, just not sure what to do with the sec^3 t

OpenStudy (anonymous):

ah, okay.

OpenStudy (turingtest):

^there it is this will be one of those integrals that repeats, so you will wind up adding the integral of sec^3 to both sides you will still have to integrate sec though, hope you remember how to do that ;)

OpenStudy (anonymous):

I do :)

OpenStudy (anonymous):

correst